In the past year, Pembina Company had total revenue of $2,540,000, cost of goods sold of $1,016,000 (before adjustment for over- or underapplied overhead), administrative expenses of $590,000, and selling expenses of $360,000. During the year, overhead was applied using a predetermined rate of 62 percent of direct labor cost. Actual direct labor was $640,000. Actual overhead was $318,800. The ending balances in the inventory accounts (prior to adjustment for overapplied overhead) are:

Raw Materials Inventory $33,000
Work in Process Inventory 69,000
Finished Goods Inventory 55,000


A. Calculate net income treating the amount of overrapplied overhead as immaterial and assigning it to Cost of Goods Sold.
B. Calculate net income treating the amount of overapplied overhead as material and apportioning it to the appropriate inventory accounts and Cost of Goods Sold. (Round proportions to 3 decimal places, e.g. 15.253 and final answer to the nearest whole dollar, e.g. 5,275.)
Answer #1

a) Overhead applied = 640000*62% = $396800

Actual overhead = 318800

Overapplied overhead = 396800-318800 = $78000

Adjusted cost of goods sold = 1016000-78000 = $938000

Net income = 2540000-938000-590000-360000 = $652000

b) overapplied overhead = 78000

Overapplied overhead allocated to cost of goods sold :78000/1140000*1016000 = 69516

Adjusted cost of goods sold = 1016000-69516 = $946484

Net income = 2540000-946484-590000-360000 = $643516